### v1

@@ 1.3 Main contributions and paper outline
The paper is organized as follows. Section 2 provides a common background knowledge of gels and the basic formulations of the chosen theories using a unified notation. Section 3 is devoted to the mathematical classification and numerical approximation using the FEM in variational settings. Section 4 explains the general setting of the numerical simulations campaign. Section 5 presents the simulation results for some well-known prototype problems of the diffusion-deformation of hydrogels. Concluding remarks are given in Section 6.
@@ 2.4 Constitutive equations: fluid flux
[…]
@@ 2.5.2 Compressible models
Constitutive model III: The following formulation is based on the ideas by [7]. To be consistent with the original model, the swelling volume change $J_{f}$ is considered in place of $c_{R}$ within the formulation, but we highlight that this is straight linked through equation (5). Recalling that $J=\text{det}({\bf F})$, the first Piola-Kirchhoff stress tensor considers the entropic and energetic components in equations (32) and (39)${}_{1}$, reading from equation (15):
@@
The chemical potential is obtained from equations (16), (29), (30) and (39)${}_{1}$ as:
@@
Constitutive models IV and V: Starting from the original incompressible model introduced in [12] (constitutive model II), the same group of authors presented alternative compressible formulations by adding energetic components as given in equations (39)${}_{2}$ ([13], constitutive model IV) and (${}_{3}$ ([14], constitutive model V). To be consistent with the original models, the polymer volume fraction $\phi$ is considered in place of $c_{R}$ within the formulation, but we highlight that this is straight linked through equation (6). Hence, recalling that $J=\text{det}({\bf F})$, the first Piola-Kirchhoff stress tensor follows, for constitutive model IV, from equation (15) with equations (32) and (39)${}_{2}$ as:
@@
and, for constitutive model V with (${}_{3}$, as:
@@
Equation (16), together with equation (30), (39)${}_{2}$ and (${}_{3}$, yields the chemical potential:
@@ 2.5.3 Preliminary comparisons between models and final considerations
In the context of compressible models, we can gain valuable insights by examining how different choices of the energy component $\Psi_{R}^{\text{en}}$ impact the stress constitutive relationships. To illustrate this, let us focus on equation (40) within constitutive model III, which primarily penalizes substantial elastic deformations, that is, when $|J-J_{f}|=|J_{e}-1|J_{f}\gg 0$. In contrast, constitutive models IV and V in equations (2.5.2) incorporate penalties for both significant elastic deformations and the fully swollen state, as evidenced by the behavior of $|\ln(J\phi)|=|\ln(J_{e})|$, which approaches +$\infty$+ ∞ when $|J_{e}-1|\gg 0$, and inversely, it approaches $\infty$- ∞ as the polymer volume fraction $\phi$ approaches 00. However, it is worth noting that constitutive models IV and V diverge from each other in treating large swelling deformations. Specifically, constitutive model V penalizes these deformations, occurring when $\phi^{-1}=J_{f}\rightarrow+\infty$, whereas constitutive model IV does not.
Furthermore, as previously highlighted, constitutive models IV and V serve as the compressible counterparts to constitutive model II. This connection becomes evident when we observe that the Lagrange multiplier $P$ in equation (37) is effectively replaced by a term related to bulk modulus in equations (2.5.2). However, from a numerical implementation standpoint, this substitution carries significant implications. In general, $P$ represents an additional primary variable that must be determined through the stationary conditions of the constrained functional (17) with respect to the Lagrange multiplier. Only under specific circumstances, such as a traction-free condition in the presence of plane stresses, can $P$ be directly defined from equilibrium conditions. In these instances, constitutive model II simplifies to having only two primary variables, namely, u and $\phi$ (as demonstrated in, for example, equation (78) in the subsequent Section 5.2). From this point forward, we will exclusively focus on these special cases.
@@ 2.6 Weak formulations
with ``:"normal-:normal-`normal-`normal-"``~{}:~{}"` ` : " denoting the double contraction of the second-order tensors $\mathbf{P}$ and $\nabla\mathbf{v}$, where $\mathbf{P}(\mathbf{u},\varphi)$ is defined by either equation (34), (37), (40), or (2.5.2), depending of the constitutive model adopted. Whereas ${\bf M}({\bf u},\varphi)$ is given by equation (44).
@@ 3.4 Numerical solution
with the joint unknown, trial function, $\mathbf{U}_{h}^{n}:=(\mathbf{u}_{h}^{n},\varphi_{h}^{n})$ and the joint test function $\mathbf{\Psi}_{h}:=(\mathbf{v}_{h},q_{h})$. This corresponds to Step 4 [45][Section 7.8.4.1]. Then, we can proceed with Step 5 (Newton’s method) in [45][Section 7.8.4.1] for the nonlinear solution. The resulting solving system is analogous to equations (3.4) for Formulation 5, but with $\mathbf{U}_{h}^{n}$ replaced by $\hat{\mathbf{U}}_{h}^{n,j}:=(\hat{\bf u}_{h}^{n,j},\hat{\varphi}_{h}^{n,j})$ that collects trial values of unknown fields at step $j$ within the iterative solution scheme.
@@
Finally, it is well known that Dirichlet boundary conditions on the chemical potential, see equation (11)${}_{4}$, might be the source of spurious numerical oscillations due to large sudden pressures within the hydrogel at the start of the simulation. Hence, whenever needed, such boundary condition is incrementally applied during the simulation, a strategy known as time-ramping boundary condition. In each case, the boundary condition is increased as fast as possible to meet a good compromise between numerical stability and physical reality. This is achieved by introducing a time-dependent exponential term that multiplies the Dirichlet boundary condition, i.e., $h(\varphi,t)=\bar{h}(\varphi)\left(1-\text{exp}(-\alpha_{r}t)\right)$, where $\alpha_{r}$ is a positive constant that determines the rate of ramping.
@@ 4 General settings for the numerical simulation campaign
Two campaigns of numerical simulations will be presented. Section 5 addresses different representative prototype problems by adopting parameter settings presented in the original papers where each constitutive model has been originally presented. The aim is to explore the robustness of the developed numerical implementations. Section 6 presents a unified benchmark problem to have a unique reference simulation example that shows the differences in the response for each constitutive equation.
@@ 5.2 Constitutive model II
In [14], the coupled problem was solved using the commercial software Abaqus, via an UMAT routine. The coupled problem was solved for $\phi$, $\mu$, and $\mathbf{u}$. In particular, $\phi$ was defined as a local variable, and the nonlinear equation (38) was solved at each Gauss integration point. This approach allowed to fully determine the time derivative in equation (48)${}_{2}$ at each time instance. Thus, a static problem for $\mu$ and $\mathbf{u}$ is solved also at each time step. A similar idea was adopted by [7].
@@ 5.4 Constitutive models IV and V
The fourth and fifth models under consideration are composed of the balance equations (10) and (11) together with constitutive equations (2.5.2) and (46) for the stress and chemical potential, respectively.

### v2

@@ 1.3 Main contributions and paper outline
The paper is organized as follows. Section 2 provides a common background knowledge of gels and the basic formulations of the chosen theories using a unified notation. Section 3 is devoted to the mathematical classification and numerical approximation using the FEM in variational settings. Section 4 explains the general setting of the numerical simulations campaign. Section 5 presents the simulation results for some well-known prototype problems of the diffusion-deformation of hydrogels. Section 6 presents a unified benchmark problem for the diffusion-deformation of hydrogels. Concluding remarks are given in Section 7.
@@ 2.4 Constitutive equations: fluid flux
Notice that to make the units of equation (28) consistent, $c_{R}$ must have units of [mol], $k_{B}$ units of [J K-1], $T$ units of [K], $\mu$ units of [J mol-1], $D$ units of [m2 s-1], and $\bm{b}$ is dimensionless. Some authors have defined constitutive equation (26) in terms of $RT$, where $R$ is the gasses constant, (see, e.g., [13] or [14]). In this case, $\mu$ units of [J].
@@ 2.5.2 Compressible models
Constitutive model III: The following formulation is based on the ideas by [7]. To be consistent with the original model, the swelling volume change $J_{f}$ is considered in place of $c_{R}$ within the formulation, but we highlight that this is straight linked through equation (5). Recalling that $J=\text{det}({\bf F})$, the first Piola-Kirchhoff stress tensor considers the entropic and energetic components in equations (32) and (39)1, reading from equation (15):
@@
The chemical potential is obtained from equations (16), (29), (30) and (39)1 as:
@@
Constitutive models IV and V: Starting from the original incompressible model introduced in [12] (constitutive model II), the same group of authors presented alternative compressible formulations by adding energetic components as given in equations (39)2 ([13], constitutive model IV) and (39)3 ([14], constitutive model V). To be consistent with the original models, the polymer volume fraction $\phi$ is considered in place of $c_{R}$ within the formulation, but we highlight that this is straight linked through equation (6). Hence, recalling that $J=\text{det}({\bf F})$, the first Piola-Kirchhoff stress tensor follows, for constitutive model IV, from equation (15) with equations (32) and (39)2 as:
@@
and, for constitutive model V with (39)3, as:
@@
Equation (16), together with equation (30), (39)2 and (39)3, yields the chemical potential:
@@ 2.5.3 Preliminary comparisons between models and final considerations
In the context of compressible models, we can gain valuable insights by examining how different choices of the energy component $\Psi_{R}^{\text{en}}$ impact the stress constitutive relationships. To illustrate this, let us focus on equation (40) within constitutive model III, which primarily penalizes substantial elastic deformations, that is, when $|J-J_{f}|=|J_{e}-1|J_{f}\gg 0$. In contrast, constitutive models IV and V in equations (45) incorporate penalties for both significant elastic deformations and the fully swollen state, as evidenced by the behavior of $|\ln(J\phi)|=|\ln(J_{e})|$, which approaches +$\infty$+ ∞ when $|J_{e}-1|\gg 0$, and inversely, it approaches $\infty$- ∞ as the polymer volume fraction $\phi$ approaches 00. However, it is worth noting that constitutive models IV and V diverge from each other in treating large swelling deformations. Specifically, constitutive model V penalizes these deformations, occurring when $\phi^{-1}=J_{f}\rightarrow+\infty$, whereas constitutive model IV does not.
Furthermore, as previously highlighted, constitutive models IV and V serve as the compressible counterparts to constitutive model II. This connection becomes evident when we observe that the Lagrange multiplier $P$ in equation (37) is effectively replaced by a term related to bulk modulus in equations (45). However, from a numerical implementation standpoint, this substitution carries significant implications. In general, $P$ represents an additional primary variable that must be determined through the stationary conditions of the constrained functional (17) with respect to the Lagrange multiplier. Only under specific circumstances, such as a traction-free condition in the presence of plane stresses, can $P$ be directly defined from equilibrium conditions. In these instances, constitutive model II simplifies to having only two primary variables, namely, u and $\phi$ (as demonstrated in, for example, equation (78) in the subsequent Section 5.2). From this point forward, we will exclusively focus on these special cases.
@@ 2.6 Weak formulations
with ``:":``"``~{}:~{}"` ` : " denoting the double contraction of the second-order tensors $\mathbf{P}$ and $\nabla\mathbf{v}$, where $\mathbf{P}(\mathbf{u},\varphi)$ is defined by either equation (34), (37), (40), or (45), depending of the constitutive model adopted. Whereas ${\bf M}({\bf u},\varphi)$ is given by equation (44).
@@ 3.4 Numerical solution
with the joint unknown, trial function, $\mathbf{U}_{h}^{n}:=(\mathbf{u}_{h}^{n},\varphi_{h}^{n})$ and the joint test function $\mathbf{\Psi}_{h}:=(\mathbf{v}_{h},q_{h})$. This corresponds to Step 4 [45][Section 7.8.4.1]. Then, we can proceed with Step 5 (Newton’s method) in [45][Section 7.8.4.1] for the nonlinear solution. The resulting solving system is analogous to equations (67) for Formulation 5, but with $\mathbf{U}_{h}^{n}$ replaced by $\hat{\mathbf{U}}_{h}^{n,j}:=(\hat{\bf u}_{h}^{n,j},\hat{\varphi}_{h}^{n,j})$ that collects trial values of unknown fields at step $j$ within the iterative solution scheme.
@@
Finally, it is well known that Dirichlet boundary conditions on the chemical potential, see equation (11)4, might be the source of spurious numerical oscillations due to large sudden pressures within the hydrogel at the start of the simulation. Hence, whenever needed, such boundary condition is incrementally applied during the simulation, a strategy known as time-ramping boundary condition. In each case, the boundary condition is increased as fast as possible to meet a good compromise between numerical stability and physical reality. This is achieved by introducing a time-dependent exponential term that multiplies the Dirichlet boundary condition, i.e., $h(\varphi,t)=\bar{h}(\varphi)\left(1-\text{exp}(-\alpha_{r}t)\right)$, where $\alpha_{r}$ is a positive constant that determines the rate of ramping.
@@ 4 General settings for the numerical simulation campaign
Two campaigns of numerical simulations will be presented. Section 5 addresses different representative prototype problems by adopting parameter settings presented in the original papers where each constitutive model has been originally presented. The aim is to explore the robustness of the developed numerical implementations. Section LABEL:res:benchmark presents a unified benchmark problem to have a unique reference simulation example that shows the differences in the response for each constitutive equation.
@@ 5.2 Constitutive model II
In [14], the coupled problem was solved using the commercial software Abaqus, via an UMAT routine. The coupled problem was solved for $\phi$, $\mu$, and $\mathbf{u}$. In particular, $\phi$ was defined as a local variable, and the nonlinear equation (38) was solved at each Gauss integration point. This approach allowed to fully determine the time derivative in equation (48)2 at each time instance. Thus, a static problem for $\mu$ and $\mathbf{u}$ is solved also at each time step. A similar idea was adopted by [7].
@@ 5.4 Constitutive models IV and V
The fourth and fifth models under consideration are composed of the balance equations (10) and (11) together with constitutive equations (45) and (46) for the stress and chemical potential, respectively.